OpenStudy (anonymous):

What is the y-coordinate of the ordered pair that satisfies the system of linear equations? A. 2 B. 1 C. -1 D. -2

OpenStudy (calculusxy):

what linear equation?

OpenStudy (anonymous):

@calculusxy I don't even know, that's the whole equation.

OpenStudy (calculusxy):

i am confused with this.

OpenStudy (anonymous):

@calcalusxy Oh! It's this: 6x + 4y = -14 -x - 5y = 11 Sorry, didn't know what you were talking about for a second!

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@calculusxy if you don't know tell me